OpenStudy (swag):

x + 3 = x + 3 This is an example of which property below? Reflexive Property Commutative Property for Addition Addition Property of Equality Transitive Property of Equality Multiplication Property of Equality

OpenStudy (hunus):

That's the reflexive property.

OpenStudy (swag):

Are you sure

OpenStudy (hunus):

Yes. The Reflexive property tells us that a=a (x+3)=(x+3)

OpenStudy (swag):

alright thank you

OpenStudy (hunus):

Yup